A horizontal rectangular surface has dimensions 2.50 cm by 3.00 cm and is in a uniform magnetic field that is directed at an angle of 34.5 degree above the horizontal.
What must the magnitude of the magnetic field be in order to produce a flux of 4.50
Answer #1

Flux = B.dS

B= ?

dS = 2.5 X 3.5 cm2

Angle = 90 - 34.5 = 55.5

B = 4.50×10−4 / (0.025 X 0.035 X cos 55.5 ) = 0.908 T